OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 - 16 = 0 and x2 = -2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (anonymous):

Part 1 x² - 16 = 0 => (x+4)(x-4)=0 => x = +4 or x = -4. x² = -2x + 24 => x² +2x - 24=0 => (x+6)(x-4)=0 => x = 4 or x = -6

OpenStudy (anonymous):

thank you